1. The nurse in the infusion clinic is preparing to give DOXOrubicin 100 mg IV to a 23 year old with leukemia who weighs 47 kg. The drug reference manual states the following about DOXOrubicin:

Route/Dosage
DOXOrubicin
IV (Adults): 60 – 75 mg / m2 / daily, repeat q.21 days.

The nurse finds in the electronic medical record that the patient’s BSA is 1.51 m2. Is the ordered dose safe for this patient? __________

2. The physician orders gabapentin 200 mg t.i.d. orally for an 8 year old. The patient weighs 88 lb. The reference manual states: PO (Children 5 – 12 years): 10 – 15 mg / kg / day in three divided doses initially. What is the maximum safe initial recommended dose for this child? _________ mg / dose

3. Order: Acetaminophen elixir 10 mg/kg po.

     Child weight is 10 pounds.

     Available: Acetaminophen elixir 160 mg/ 5ml.

     How much Acetaminophen will you administer?
Answer #1

ANSWER

1.Prepared dose of Doxorubicin for administration = 100 mg.

    Drug reference range for adults = 60 - 75 mg / m2.

               BSA of the patient = 1.51 m2.

   Safe dose for the patient

                     Based on drug reference mannual

                                       60 - 75 mg daily for patient with 1m2 body surface area.

                                       So dose can given for a patient with 1.51m2 is

                                                         = 60 * 1.51 to 75 * 1.51

                                                        = 90.6 mg to 113.25 mg.

     * Ordered dose is 100 mg which is between the normal reference range for doxorubicin

        so it is safe for the patient.

2. Ordered dose of drug as per physician's order = 200mg.(t.i.d)

              Weight of the patient = 88 lb (40 kg)

                Reference mannual range = 10 - 15 mg /kg / day ( 3 divided doses).

    Maximum safe initial recommended dose

                    Total dose can given for the child with body weight 88 lb(40 kg) is

                                                                   = 10 * 40 to 15 * 40

                                                                   = 400 mg to 600 mg.

                     Initially three divided doses = 400 / 3 to 600 / 3    

   = 133.3mg to 200mg.

   So maximum safe initial recommended dose is 200 mg.

3. Amount of drug acetaminophen administered.

   Formula

Amount administered = Amount desired (D) / Amount in hand(H) * Quantity(Q)

   Amount desired = 10mg / kg.

child weight = 10 pound ( 4.5 kg)

= 10 mg * 4.5 kg

       D = 45 mg.   

   Amount in hand = 160 mg.

   Quantity   = 5 ml.

               Amount administered = D / H * Q

                                                         = 45 / 160 *5

                                                         = 0.28 * 5

                                                        = 1.4 ml.